a body has a mas 80 kilogram and height 250 km. find its weight using universal law of gravitation in newton and in pounds??
1
Expert's answer
2012-10-25T11:26:07-0400
According to the Newton's law of universal gravitation we have:

F = int(...)dh (from A to B)

F = int ( 6.67384×10^(-11)[m³/kg/s²] * 5.97219×10^24[kg] * 80[kg]/250000[m] / h² )dh (from 6353000[m] to 6353000[m]+250000[m]) ≈ 760.1149[N] = 760.1149[kg·m/s²] = 760.1149/0.45359237[pounds·m/s²] = 1675.7665[pounds·m/s²].